Find the eigen value and eigen function of homogeneous integral equation

y(x)=
1
Expert's answer
2021-09-06T16:05:35-0400

This question is incomplete but we can take "y(x)=\lambda \int_{0}^{1} e^{x+1} y(t) d t" to understand the problem.

The given equation can be written as

"y(x)=\lambda e^{x} \int_{0}^{1} e^{t} y(t) d t \ldots(\mathrm{i})"

Assume that "\quad C=\int_{0}^{1} e^{t} y(t) d t \ldots(\mathrm{ii})"

Then, from Eq. (i), we have "u(x)=\lambda C e^{x} \ldots . . (iii)"

"\Rightarrow y(t)=\lambda C e^{t}"

Putting this value in eq. (ii), we get

"\begin{aligned}
&C=\int_{0}^{1} e^{t}\left(\lambda C e^{t}\right) d t=\lambda C\left[\frac{e^{2 t}}{2}\right]_{0}^{1} \\
&\Rightarrow C=\frac{\lambda C}{2}\left(e^{2}-1\right) \Rightarrow C\left[1-\frac{\lambda}{2}\right]\left(e^{2}-1\right)=0
\end{aligned}"

If "\mathrm{C}=0" , then Eq. (iii) gives "{y}({x})=0" .

We, therefore assume that for non-zero solution of Eq. (i).

If "C \neq 0" , then Eq. (iii) gives

"1-\frac{\lambda}{2}\left(e^{2}-1\right)=0"

"\Rightarrow \lambda=\frac{2}{e^{2}-1}"

Which is an eigen value of Eq. (i).

To find the corresponding eigen function, putting the value of "\lambda" in Eq. (iii), we get

"y(x)=\frac{2 C}{e^{2}-1} \cdot e^{x}"

The eigen function is "\mathrm{e}^{{x}}."